OpenStudy (anonymous):

Help? Eliminate the parameter. x = 4 cos t, y = 4 sin t

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

Alternatively, you can take the fact that the parametric equations describe a circle with radius 4, so you have (4cost)2+(4sint)2=42x2+y2=16

OpenStudy (anonymous):

Okay, wait... Can you do that in the equation thing? I can't tell what exactly you did.

OpenStudy (anonymous):

I know we have to square both sides first..

OpenStudy (anonymous):

@sexyperson

OpenStudy (anonymous):

Where did you get the 42?

OpenStudy (anonymous):

look at this example from openstudy

OpenStudy (anonymous):

I still don't get it..

OpenStudy (anonymous):

Is that supposed to be (4cost)^2+(4sint)^2=4^2+x^2+y^2=16?

OpenStudy (anonymous):

How do you know it equals 16?

OpenStudy (anonymous):

@Light&Happiness

OpenStudy (anonymous):

If I square both 4's inside the parenthesis, I don't get 16, I get 32..???

OpenStudy (whpalmer4):

Okay, I don't do parametric equations very often, haven't for many years, but yes, this describes a circle with radius 4. Here's a plot of the two parametric equations: remember that one is the x coordinate, and the other is the y coordinate...

OpenStudy (whpalmer4):

So you can see that as \(t\) goes from \(0\) to \(2\pi\), the y coordinate goes from 4 to 0 to -4 and back to 0 and then to 4, and the x coordinate simultaneously goes from 0 to 4 to 0 to -4 to 0

OpenStudy (whpalmer4):

If you look at the graph of a circle with radius 4, that's what happens as you start at the top and go clockwise around the circle, right?

OpenStudy (anonymous):

Yes..?

OpenStudy (whpalmer4):

Here it is showing the angle from vertical (in degrees) as the x coordinate.

OpenStudy (anonymous):

I'm confused on this part:(4cost)2+(4sint)2=42x2+y2=16

OpenStudy (anonymous):

:/

OpenStudy (whpalmer4):

then ignore it, it looks wrong to me.

OpenStudy (whpalmer4):

A circle with radius 4 with center at the origin will have what equation? \[(x-h)^2 + (y-k)^2 = r^2\]

OpenStudy (anonymous):

I thought so.. So we need to use that? Do we need to divide by the 4's?

OpenStudy (whpalmer4):

what would your equation for a circle with radius 4 centered at the origin be?

OpenStudy (whpalmer4):

formula gives you a circle with radius \(r\) centered at \((h,k)\). We want center at \((0,0)\) so \(h = 0, k = 0\) and radius \(4\) so \(r = 4\): \[(x-0)^2 + (y-0)^2 = 4^2\] or\[x^2+y^2=16\] Yes, I realize that I have not eliminated the parameter in the way they want you to do. I make no pretense about doing so — I'm trying to get you to learn how to reject bogus information.

OpenStudy (anonymous):

(x-0)^2+(y-0)^2=4^2?

OpenStudy (anonymous):

Okay, whatever works. :)

OpenStudy (whpalmer4):

Yes. Now, if you know what you're supposed to be getting, you have a better shot at evaluating whether the method someone is telling you how to do this is going to work, right?

OpenStudy (anonymous):

I guess so... But that's my answer?... x^2+y^2=16. What happened to my t? I'm lost..

OpenStudy (whpalmer4):

you're eliminating the parametric variable, right? t is the parametric variable. So yes, x^2+y^2=16 is the answer here, but to the extent you're supposed to show your work and know how to do this, you're not exactly done yet :-)

OpenStudy (whpalmer4):

Here, if you read this webpage, partway down the page they do an equivalent problem: http://colalg.math.csusb.edu/~devel/precalcdemo/param/src/param.html

OpenStudy (anonymous):

Thanks, I'll check it out! God bless you! :D